The EUR/USD spot quote on the interbank market is 1.3140/43. Euroland interest rates are 2.065/2.135% and US interest rates are 1.424/1.598%. Using this information and round-trip transactions, calculate the range of bid-offer quotes that ensure no arbitrage opportunities exist over a 3-month (92-day) period. Explain how you arrived at your answer. (For calculation purposes, assume 360 days in the financial year).The EUR/USD spot quote on the interbank market is 1.3140/43. Euroland interest rates are 2.065/2.135% and US interest rates are 1.424/1.598%. Using this information and round-trip transactions, calculate the range of bid-offer quotes that ensure no arbitrage opportunities exist over a 3-month (92-day) period. Explain how you arrived at your answer. (For calculation purposes, assume 360 days in the financial year).
Answer #1

As per interest rate parity :it is a theory of interest rate differential between two countries is equal to the difference between forward rates and spot exchange rates Forward rate=Spotrate(+rate of home currency) /(1+rate of foreign currency) .1 Eur=1.3140$ so 1.3140*(1+.01424)/(1+.02065)= 1.3141*.99372=1.30585 .Hence forward rate is 1.30585. Rate of premium / discount = Forward -Spot/Spot*360/n*100. 1.30585-1.314=-.00815 /1.3140*360/92*100= -2.42704 percent. Now we calculate dame way for ask rate 1 Euro =1.3143 so 1.3143*(1+0.01598)/(1+.02135)=1.353594 so Forward rate-Spot rate /Spot*360/n*100=1.35359-1.3143/1.3143*360/92*100= 11.697746